# Structured Representation

### v1

@@ Abstract
Invariant representations are core to representation learning, yet a central challenge remains: uncovering invariants that are stable and transferable without suppressing task-relevant signals. This raises fundamental questions, requiring further inquiry, about the appropriate level of abstraction at which such invariants should be defined, and which aspects of a system they should characterize. Interpretation of the environment relies on abstract knowledge structures to make sense of the current state, which leads to interactions, essential drivers of learning and knowledge acquisition. We posit that interpretation operates at the level of higher-order relational knowledge; hence, invariant structures must be where knowledge resides, specifically, as partitions defined by the closure of relational paths within an abstract knowledge space. These partitions serve as the core invariant representations, forming the structural substrate where knowledge is stored and learning occurs. On the other hand, inter-partition connectors enable the deployment of these knowledge partitions encoding task-relevant transitions. Thus, invariant partitions provide the foundational primitives of structured representation. We formalize the computational foundations for structured representation of the invariant partitions based on closed semiring, a relational algebraic structure.
@@ 1 Introduction
(b) Structured representation
Figure 1: a) Interpretation entails posing internal questions. For example, when examining an image of the historical artifact, The Antikythera mechanism Joy of Museum (2018), one’s interpretation is shaped by the relational links drawn from prior knowledge. What invariant structures are fundamental to learning relational knowledge representations? This underscores an inquiry in structured representations. b) Structured representation, where invariant representations are at a higher level of abstract knowledge space defined by partition blocks and the task relevant connectors, provide a promising offer to address core challenges in contemporary representation learning.
The motivation for this problem arises from the process of interpretation. Consider the image in Figure 1, which depicts a historical artifact, and contemplate how different individuals might interpret it. Their interpretations vary depending on prior knowledge — some may identify familiar components, while others may be drawn to questions about its composition, origin, lighting effects, or functional purpose. Is it a steering unit, a computational device, or something entirely different? Even in everyday tasks, people interpret and contextualize the world before interacting with objects in it.These varying perspectives do not simply reflect differences in knowledge, but rather alternative views on a common subject. Yet, a common thread emerges: people strive for closure in understanding via interpretation grounded in internal abstract knowledge systems. This process drives abstract reasoning, supports learning and knowledge acquisition, and guides meaningful interactions with the world. We argue that the pursuit of closure should guide the development of representations that bridge internal knowledge structures with interactions. Unlike arbitrarily imposed forms, closure-driven representations are grounded in interpretations. This perspective prompts several foundational questions: What are the goals, mechanisms, and representational structures that mediate an agent’s internal knowledge and its interactions? These inquiries are central to advancing our understanding of representation and learning.
@@
As a result, we posit that invariant representations in natural systems are structured partitions in an abstract knowledge space, formalized through the principle of closure. Structured representations must satisfy three core requirements: Chaining refers to the ability to sequentially compose relations into composite relations within partitions and make connections between partitions. Choice entails selecting among multiple viable alternatives through querying and task-dependent filtering. Closure ensures and checks consistency within partitions, enforces conditions for relational chaining, and enables redirection to alternative paths when necessary. Basis paths thus emerge as the fundamental invariant representations. The path-based invariant representation view focuses on what topological representations are changing, and this is where learning takes place. These partitions function as structural knowledge units, each with defined entry and exit points, enabling seamless reasoning within the partition. The disjoint partitions can then be flexibly composed through connectors to support task-specific reasoning. This perspective simultaneously addresses the dual needs of stable representations and task-relevant flexibility. Intuitively, it is analogous to how people can reason in structured blocks: reasoning proceeds effortlessly within a partition block, while problem solving revolves around how to stitch together multiple such blocks. Therefore, structured relational computations must be encapsulated within the primitives of structured representations—the three C’s of structured representations: closure, choice and chaining. In this article, we formalize the computational foundations for structured representations using a relational algebraic structure closed semiring, which satisfies the essential properties required for representing and reasoning about relational structures in natural systems.
@@
Learning invariant representations is a foundational objective in machine learning. The aim is to map raw observations into compact representations that remain stable under input variations while retaining predictive utility across contexts. A broad range of approaches - such as representation learning frameworks Bengio et al. (2013), contrastive learning Chen et al. (2020), and domain-adversarial training Ganin et al. (2016)— seek to encode invariances either explicitly or implicitly by embedding them within objective functions. However, a persistent challenge lies in balancing invariance with discriminability: representations that are invariant to spurious variations may also suppress task-relevant signals Locatello et al. (2019); Ahuja et al. (2020). Domain generalization aims to address this challenge by training models that generalize effectively to unseen target domains Wang et al. (2022). Strategies in this space include learning domain-invariant features to align source and target distributions Muandet et al. (2013); Li et al. (2018), or directly mitigating domain-specific biases. However, inability to anticipate the nature of distributional shifts in target domains often results in models that overfit to the source domains and fail to capture the essential transferable structure Gulrajani and Lopez-Paz (2020). The premise of domain generalization thus hinges on the assumptions about source-target relationships and the ability of models to acquire representations that are truly transferable.
Relational reasoning can be categorized into three core types: induction, deduction, and transduction. Induction refers to generalizing from specific observed instances to broader patterns - supervised learning. Deduction involves deriving logical conclusions from general premises — logic or symbolic systems. Transduction infers new examples directly from known examples without explicitly learning a general rule — semi-supervised methods and graph neural networks. Recent developments in neural architectures, including graph neural networks Gilmer et al. (2017) and transformers Vaswani et al. (2017), show promise in relational reasoning by modeling interactions among entities Battaglia et al. (2018) or inducing symbolic abstractions Evans et al. (2018). Nevertheless, these models lack mechanisms to represent core relational operations - composition, coherence, inversion - limiting their ability in complex reasoning tasks.
@@ 2 Relational Algebraic Structures
Although tensor computations can be used to implement relational algebraic operations, several aspects of critical reasoning in relation computations require a highly structured implementation approach. In particular, relations can be viewed as operators, hypergraphs, or k-ary typed partial functions that have algebraic composition and transform requirements. Our implementation closely adheres to a tensor approach while enforcing these requirements. We treat relations as typed partial functions, using subscripts to indicate inputs and superscripts to indicate outputs. Thus, relations are reduced to simpler tensors by currying (partial evaluation), and using this strategy, we can represent a network of validly composed relations as a directed graph over curried relations. By structuring all inputs and outputs as (type, data) tuples, $[R_{t_{i_{1}},x_{i_{1}}}^{t_{o_{1}},y_{o_{1}}},R_{t_{i_{2}},x_{i_{2}}}^{t_{o_{2}},y_{o_{2}}},\cdots,R_{t_{i_{n}},x_{i_{n}}}^{t_{o_{n}},y_{o_{n}}}]$, relational compatibility can be decided locally by matching tensor index.
Semirings provide a convenient computational framework for these operations Dolan (2013); Gondran and Minoux (2008). They are algebraic structures that have distributive addition and multiplication without the need for multiplicative inverses. Semirings ubiquitously occur in problems with two kinds of composition: parallel (choice) and serial (chaining). Closed semirings have an additional operation of closure, an important algebraic operation in ensuring coherence structure in natural systems.
@@
  • 00 is absorbing for $\otimes$: $00\otimes a=a\otimes 0=$ 0
A closed semiring is a semiring with an additional unary operation called closure $(S,\oplus,\otimes,{}^{*},0,1)$ which satisfies $a^{*}=1\oplus(a\otimes a^{*})=1\oplus(a^{*}\otimes a)$ for all $a\in S$. $a^{*}$ is intuitively the union of all chainings of $a$. The closure operator generalizes the concept of reflexive-transitive closure in relational semirings or path closure in graphs.
Relational closure A directed graph $G=(V,E)$ is given by a set of nodes $V$ and a set of edges $E$. A relation $r$ is an ordered pair $(u,v)$ denoted by $urv$, where $u,v\in V$. The graph has a finite number of nodes, $n=|V|$, and a relation matrix $R$. Ownership relations are common in the natural world, and binary relational graphs can be represented by binary relational matrices Givant and Givant (2017); Doumane (2021) as $R[i,j]=1\text{ if }irj,0\text{ otherwise}$. If $u$ and $v$ are two nodes in a graph, then a path between them goes from $u$ to $v$ along the graph’s edges, i.e. a path is composed of relations between the two nodes. If $urv$ and $vrw$, then $urw$, implying that the relation is transitive. Adding relation matrices is to take the union of edges, and multiplication of relation matrices is relational chaining. The composition operation enables us to create new composite relations from existing relations. As the graph has a finite number of nodes, the transitive closure of its relation matrix is the union of the first $n$ powers of R, $R^{*}=R\cup R^{2}\cup...\cup R^{n}=\cup_{i=1}^{n}R^{i}$. This reachability relation informs if a node $j$ is reachable from another node $i$ through any number of edges, therefore, it is relational closure. Reflexive transitive closure can be computed using Floyd (1962). Lehmann (1977) shows that the closure matrix can be broken down into block structures that are recursively computable Kozen (2012). These findings are significant because closure in binary matrices assures that relational reasoning or answer-finding have a coherent chaining structure and a termination point.
@@ 3 Structured relational computation
Figure 2: Overview of the core elements of the structured representations. Directed graphs are available after partial evaluation, as discussed above. a) Abstract knowledge is in the form of block partitions. In relational systems, its not entity level relations which are property to property relations. b) Relations can be chained together to form composite relations within block partitions, which define encapsulated units of abstraction and have an entry and exit points. c) Partitions can be queries for choices through forward or inverse query. Forward query results in which properties are related to inquiry while inverse query results in from which properties the answer can be arrived at. d) Closure illustrates reflexive transitive closure and block partition structures in a simple example. These three element together conform to the working of coherent and compositional systems.
@@
Given a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{P},\mathcal{E})$, the partially ordered set (poset) $(\mathcal{P},\preceq)$ is defined by the directed relations $urv$ where the relation connects node $u$ to $v$ in the graph and $v$ is relational successor of $u$, $u\preceq v$ in the graph. Given a finite poset $(\mathcal{P},\preceq)$ of elements $\{x_{1},x_{2},\dots,x_{n}\}$, the zeta function $\zeta$ is defined as $\zeta(x,y)\texttt{=}1\text{ if }x\preceq y\text{ and }0\text{ otherwise}$ for all $x,y\in\mathcal{P}$. If $f:\mathcal{P}\rightarrow\mathbb{R}$ is a function on the poset, then the zeta transform of $f$ is given by $F(y)\texttt{=}\sum\limits_{x\preceq y}f(x)$ and in matrix notation, the zeta transform can be written as $F\texttt{=}Zf$ where $Z$ the zeta matrix is an $n\times n$ matrix $Z_{ij}\texttt{=}1\text{ if }x_{i}\preceq x_{j},0\text{ otherwise}$ where $F$ is the transformed function. The Möbius function $\mu$ of a poset $\mathcal{P}$ is defined recursively as $\mu(x,y)\texttt{=}-\sum\limits_{x\preceq z\prec y}\mu(x,z)\text{ if }x\prec y\text{ and }1\text{ if }x\texttt{=}y$. The Möbius transform recovers the original function $f$ from the transformed function and is given by $f(y)\texttt{=}\sum\limits_{x\preceq y}\mu(x,y)F(x)$ and in matrix form $f\texttt{=}MF$ where $M$ is the Möbius matrix given by $M[i,j]\texttt{=}\mu(x_{i},x_{j})$. Directed graphs define poset structure by chaining relations; thus, the zeta transform functions as a partial relational closure operator.
@@
where $\lambda$ appears on the diagonal, $1$’s appear on the super-diagonal just above the diagonal, and 00’s elsewhere. A square matrix $A^{n\times n}$ is said to be in Jordan normal form if it is block-diagonal with Jordan blocks along the diagonal:
@@
where each $J_{n_{i}}(\lambda_{i})$ is a Jordan block corresponding to eigenvalue $\lambda_{i}$. Each eigenvalue $\lambda_{i}$ may appear in multiple Jordan blocks and if there exists an invertible matrix $P$ such that $A=PJP^{-1}$, then $J$ is called a Jordan normal form of $A$ and it is a canonical representative of the similarity equivalence class of A. Each Jordan block $J_{n_{i}}(\lambda_{i})$ in $J$ corresponds to a chain of generalized eigenvectors associated with eigenvalue $\lambda_{i}$. The set of all these chains forms a generalized eigenspace which is invariant under A and nilpotent within each generalized eigenspace. Each block encodes its chain length. For Jordan block on relation matrix of a DAG, each block corresponds to a chain of nodes and each block represents paths. A Jordan block of size $k$ would mean a path of length $k-1$ in the DAG. To keep the mathematical notation consistent with familiar Jordan block notations in literature, we use $J$ to represent them, and denote partition blocks using $\mathcal{Z}$, as is common in representation learning. Both notations are used interchangeably for partition blocks throughout the text. Therefore, each block partitions the abstract space into internal blocks structures.
@@
Block structures function as the modus operandi for querying current understanding, see Figure 2c. Forward querying asks where one can reach from a given state through relational composition, while inverse querying asks from which prior states one could have arrived at the current state and through which relations and properties. This bidirectional mechanism of relational querying, forward and inverse, is analogous to forward and inverse planning. It constrains relational transitions by emphasizing that, before progressing arbitrarily to a new relation or node through interaction, one must query the underlying structural blocks. The choice monoid serves as a concrete view selection operation as long as we restrict choices to be grounded in data, allowing arbitrary selection of views constrained by available data. In contrast, query operations target specific ranges but may yield null results depending on the contents. These operations do not involve events and we can do all the relational chaining with them. The duality of analysis via relational paths and synthesis through the convergence upon abstract block structures anchors both inference and action in a coherent representational substrate.

### v2

# Structured Relational Representations

@@ Abstract
Invariant representations are core to representation learning, yet a central challenge remains: uncovering invariants that are stable and transferable without suppressing task-relevant signals. This raises fundamental questions, requiring further inquiry, about the appropriate level of abstraction at which such invariants should be defined and which aspects of a system they should characterize. Interpretation of the environment relies on abstract knowledge structures to make sense of the current state, which leads to interactions, essential drivers of learning and knowledge acquisition. Interpretation operates at the level of higher-order relational knowledge; hence, we propose that invariant structures must be where knowledge resides, specifically as partitions defined by the closure of relational paths within an abstract knowledge space. These partitions serve as the core invariant representations, forming the structural substrate where knowledge is stored and learning occurs. On the other hand, inter-partition connectors enable the deployment of these knowledge partitions encoding task-relevant transitions. Thus, invariant partitions provide the foundational primitives of structured representation. We formalize the computational foundations for structured relational representations of the invariant partitions based on closed semiring, a relational algebraic structure.
@@ 1 Introduction
(b) Representation learning
(c) Structured relational representations
Figure 1: a) Interpretation entails posing internal questions shaped by prior knowledge. For example, when interpreting an image of the Antikythera mechanism Joy of Museum (2018), individuals draw links from their prior knowledge. b) Representation learning maps input to latent representation space by optimizing appropriate objective functions, with the aim of learning invariant representations while also retaining task relevant predictive utility. However, this core issue persists. c) Structured relational representations, where invariant representations are at a higher knowledge level defined by partition blocks and the task relevant connectors, provide a promising offer to address core challenges in representation learning.
The motivation for this problem arises from the process of interpretation. Consider the image in Figure 1, which depicts a historical artifact, and contemplate how different individuals might interpret it. Their interpretations vary depending on prior knowledge — some may identify familiar components, while others may be drawn to questions about its composition, origin, lighting effects, or functional purpose. Is it a steering unit, a computational device, or something entirely different? Even in everyday tasks, people interpret and contextualize the world before interacting with objects in it. These varying perspectives do not simply reflect differences in knowledge, but rather alternative views on a common subject. Yet, a common thread emerges: People strive for closure in understanding via interpretation grounded in internal abstract knowledge systems. This process drives abstract reasoning, supports learning and knowledge acquisition, and guides meaningful interactions with the world. We argue that the pursuit of closure should guide the development of representations that bridge internal knowledge structures with interactions. Unlike arbitrarily imposed forms, representations via closure are grounded in interpretations. This perspective prompts several foundational questions: What are the goals, mechanisms, and representational structures that mediate an agent’s internal knowledge and its interactions? These inquiries are central to advancing our understanding of representation and learning.
@@
As a result, we propose that invariant representations in natural systems are structured partitions in an abstract knowledge space, formalized through the principle of closure. Structured representations must satisfy three core requirements: Chaining refers to the ability to sequentially compose relations into composite relations within partitions and make connections between partitions. Choice entails selecting among multiple viable alternatives through querying and task-dependent filtering. Closure ensures and checks consistency within partitions, enforces conditions for relational chaining, and enables redirection to alternative paths when necessary. Consistent paths within partitions thus emerge as the fundamental invariant representations. The path-based invariant representation view focuses on what topological representations are changing, and this is where learning takes place. These partitions function as structural knowledge units, each with defined entry and exit points, enabling seamless reasoning within the partition. The disjoint partitions can then be flexibly composed through connectors to support task-specific reasoning. This perspective simultaneously addresses the dual needs of stable representations and task-relevant flexibility. Intuitively, it is analogous to how people can reason in structured blocks: reasoning proceeds effortlessly within a partition block, while problem solving revolves around how to stitch together multiple such blocks.
Therefore, structured relational computations must be encapsulated within the primitives of structured representations—the three C’s of structured representations: closure, choice and chaining. In this article, we formalize the computational foundations for structured representations using a relational algebraic structure closed semiring, which satisfies the essential properties required for representing and reasoning about relational structures in natural systems.
@@
Learning invariant representations is a foundational goal in machine learning. The aim is to map observations into compact representations that remain stable under input variations while retaining predictive utility across contexts. A broad range of approaches, such as representation learning Bengio et al. (2013), contrastive learning Chen et al. (2020), and domain-adversarial training Ganin et al. (2016), seek to encode invariances either explicitly or implicitly by embedding them in objective functions. However, a persistent challenge lies in balancing invariance with discriminability: representations that are invariant to spurious variations may also suppress task-relevant signals Locatello et al. (2019); Ahuja et al. (2020). Domain generalization aims to address this challenge by training models that generalize effectively to unseen target domains Wang et al. (2022). Strategies include learning domain-invariant features to align source and target distributions Muandet et al. (2013); Li et al. (2018), or directly mitigating domain-specific biases. However, inability to anticipate the nature of distributional shifts in target domains often results in models that overfit to the source domains and fail to capture the transferable structure Gulrajani and Lopez-Paz (2020). The premise of domain generalization thus hinges on the assumptions about source-target relationships and the ability of models to acquire representations that are truly transferable.
Relational reasoning can be categorized into three core types: induction, deduction, and transduction. Induction refers to generalizing from specific observed instances to broader patterns - supervised learning. Deduction involves deriving logical conclusions from general premises — logic or symbolic systems. Transduction infers new examples directly from known examples without explicitly learning a general rule — semi-supervised methods and graph neural networks. Many studies in learning disentangled representations Esmaeili et al. (2019); Higgins et al. (2018); Locatello et al. (2019) and others have focused mainly on disentangling statistically independent axes or transforming object properties, in contrast, our aim is to develop representations for higher-order relational knowledge that facilitate interpretations in meta-level knowledge learning Kumar and Schrater (2024) in knowledge, interaction, and execution. Recent developments in neural architectures, including graph neural networks Gilmer et al. (2017) and transformers Vaswani et al. (2017), show promise in relational reasoning by modeling interactions among entities Battaglia et al. (2018) or inducing symbolic abstractions Evans et al. (2018). Nevertheless, these models lack mechanisms to represent core relational operations - composition, coherence, inversion - limiting their ability in complex reasoning tasks.
@@ 2 Relational Algebraic Structures
Although tensor computations can be used to implement relational algebraic operations, several aspects of critical reasoning in relation computations require a highly structured implementation approach. In particular, relations can be viewed as operators, hypergraphs, or k-ary typed partial functions that have algebraic composition and transform requirements. Our implementation closely adheres to a tensor approach while enforcing these requirements. We treat relations as typed partial functions, using subscripts to indicate inputs and superscripts to indicate outputs. Thus, relations are reduced to simpler tensors by currying (partial evaluation), and using this strategy, we can represent a network of validly composed relations as a directed graph over curried relations. By structuring all inputs and outputs as (type, data) tuples, $R_{t_{i_{1}},x_{i_{1}}}^{t_{o_{1}},y_{o_{1}}},R_{t_{i_{2}},x_{i_{2}}}^{t_{o_{2}},y_{o_{2}}},\cdots,R_{t_{i_{n}},x_{i_{n}}}^{t_{o_{n}},y_{o_{n}}}]$, relational compatibility can be decided locally by matching tensor index.
Semirings provide a convenient computational framework for these operations. They Dolan (2013); Gondran and Minoux (2008) are algebraic structures that have distributive addition and multiplication without the need for multiplicative inverses. Semirings ubiquitously occur in problems with two kinds of composition: parallel (choice) and serial (chaining). Closed semirings have an additional operation of closure, an important algebraic operation in ensuring coherence structure in natural systems.
@@
  • 0 is absorbing for $\otimes$: $0\otimes a=a\otimes 0=0$
A closed semiring is a semiring with an additional unary operation called closure $(S,\oplus,\otimes,{}^{*},0,1)$ which satisfies $a^{*}=1\oplus(a\otimes a^{*})=1\oplus(a^{*}\otimes a)$ for all $\in S. aa^{*}$ is intuitively the union of all chainings of $a$. The closure operator generalizes the concept of reflexive-transitive closure in relational semirings or path closure in graphs.
Relational closure A directed graph $G=(V,E)$ is given by a set of nodes $V$ and a set of edges $E$. A relation $r$ is an ordered pair $(u,v)$ denoted by urvurv, where $u,v\in V$. The graph has a finite number of nodes, $n=|V|$, and a relation matrix $R$. Ownership relations are common in the natural world, and binary relational graphs can be represented by binary relational matrices Givant and Givant (2017); Doumane (2021) as $R[i,j]=1\text{ if }irj,0\text{ otherwise}$. If $u$ and $v$ are two nodes in a graph, then a path between them goes from $u$ to $v$ along the graph’s edges, i.e. a path is composed of relations between the two nodes. If urvurv and vrwvrw, then urwurw, implying that the relation is transitive. Adding relation matrices is the union of edges, and multiplication of relation matrices is relational chaining. It enables us to create new composite relations from existing relations. As the graph has a finite number of nodes, the transitive closure of its relation matrix is the union of the first $n$ powers of R, $R^{*}=R\cup R^{2}\cup...\cup R^{n}=\cup_{i=1}^{n}R^{i}$. The reachability relation informs if a node $j$ is reachable from another node $i$ through any number of edges, therefore, it is relational closure. Reflexive transitive closure can be computed using Floyd (1962). Lehmann (1977) shows that the closure matrix can be broken down into block structures that are recursively computable Kozen (2012). These findings are significant because closure in binary matrices assures that relational reasoning has a coherent chaining structure and a termination point.
@@ 3 Structured Relational Computation
Figure 2: Overview of the core elements of the structured relational representations. Directed graphs are available after partial evaluation, as discussed above. a) Abstract knowledge is in the form of block partitions. In relational systems, it is not at object level but entity level relations, which are property to property relations. b) Relations can be chained together to form composite relations within block partitions, which define encapsulated units of abstraction and have an entry and exit points. c) Partitions can be queried for choices through forward or inverse query. Forward query results in which properties are related to inquiry while inverse query results in from which properties the answer can be arrived at. d) Closure illustrates reflexive transitive closure and block partition structures in a simple example. These three element together conform to the working of coherent and compositional systems.
@@
Given a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{P},\mathcal{E})$, the partially ordered set (poset) $(\mathcal{P},\preceq)$ is defined by the directed relations urvurv where the relation connects node $u$ to $v$ in the graph and $v$ is relational successor of $u$, $u\preceq v$ in the graph. Given a finite poset $(\mathcal{P},\preceq)$ of elements $\{x_{1},x_{2},\dots,x_{n}\}$, the zeta function $\zeta$ is defined as $\zeta(x,y)\texttt{=}1\text{ if }x\preceq y\text{ and }0\text{ otherwise}$ for all $x,y\in\mathcal{P}$. If $f:\mathcal{P}\rightarrow\mathbb{R}$ is a function on the poset, then the zeta transform of $f$ is given by $F(y)\texttt{=}\sum\limits_{x\preceq y}f(x)$ and in matrix notation, the zeta transform can be written as $F\texttt{=}Zf$ where $Z$ the zeta matrix is an $n\times n$ matrix $Z_{ij}\texttt{=}1\text{ if }x_{i}\preceq x_{j},0\text{ otherwise}$ where $F$ is the transformed function. The Möbius function $\mu$ of a poset $\mathcal{P}$ is defined recursively as $\mu(x,y)\texttt{=}-\sum\limits_{x\preceq z\prec y}\mu(x,z)\text{ if }x\prec y\text{ and }1\text{ if }x\texttt{=}y$. The Möbius transform recovers the original function $f$ from the transformed function and is given by $f(y)\texttt{=}\sum\limits_{x\preceq y}\mu(x,y)F(x)$ and in matrix form $f\texttt{=}MF$ where $M$ is the Möbius matrix given by $M[i,j]\texttt{=}\mu(x_{i},x_{j})$. Directed graphs define poset structure by chaining relations; thus, the zeta transform functions as a partial relational closure operator.
@@
where $\lambda$ appears on the diagonal, $1$’s appear on the super-diagonal just above the diagonal, and 0’s elsewhere. A square matrix $A^{n\times n}$ is said to be in Jordan normal form if it is block-diagonal with Jordan blocks along the diagonal:
@@
where each $J_{n_{i}}(\lambda_{i})$ is a Jordan block corresponding to eigenvalue $\lambda_{i}$. Each eigenvalue $\lambda_{i}$ may appear in multiple Jordan blocks and if there exists an invertible matrix $P$ such that $A=PJP^{-1}$, then $J$ is called a Jordan normal form of $A$ and it is a canonical representative of the similarity equivalence class of A. Each Jordan block $J_{n_{i}}(\lambda_{i})$ in $J$ corresponds to a chain of generalized eigenvectors associated with eigenvalue $\lambda_{i}$. The set of all these chains forms a generalized eigenspace which is invariant under A and nilpotent within each generalized eigenspace. Each block encodes its chain length. For Jordan block on relation matrix of a DAG, each block corresponds to a chain of nodes and each block represents paths. A Jordan block of size $k$ would mean a path of length k1k-1 in the DAG. To keep the mathematical notation consistent with familiar Jordan block notations in literature, we use $J$ to represent them, and denote partition blocks using $\mathcal{Z}$, as is common in representation learning. Both notations are used interchangeably for partition blocks throughout the text. Therefore, each block partitions the abstract space into internal blocks structures.
@@
Block structures function as the modus operandi for querying current understanding, see Figure 2c. Forward querying asks where one can reach from a given state through relational composition, while inverse querying asks from which prior states one could have arrived at the current state and through which relations and properties. This bidirectional mechanism of relational querying, forward and inverse, is analogous to forward and inverse planning. It constrains relational transitions by emphasizing that, before progressing arbitrarily to a new relation or node through interaction, one must query the underlying structural blocks. The choice monoid serves as a concrete view selection operation as long as we restrict choices to be grounded in data, allowing arbitrary selection of views constrained by available data. In contrast, query operations target specific ranges but may yield null results depending on the contents. These operations do not involve events and we can do all the relational chaining with them. The forward-backward mechanism or duality of analysis via relational paths and synthesis through the convergence upon abstract block structures anchors both inference and action in a coherent representational substrate.